OpenStudy (anonymous):

Brian is creating a collage on a piece of cardboard that has an area of 120r3 square centimeters. The collage is covered entirely by pieces of paper that do not overlap. Each piece has an area of the square root of r to the fifth power square centimeters. Use the given information to determine an expression for the total number of pieces of paper used. Someone please help me :)

jimthompson5910 (jim_thompson5910):

So the overall area is \(\Large 120r^3\) ? and each piece has an area of \(\Large \sqrt{r^5}\) ?

OpenStudy (anonymous):

Yess

jimthompson5910 (jim_thompson5910):

let x = number of pieces needed

jimthompson5910 (jim_thompson5910):

if you have x pieces, each of area with \(\Large \sqrt{r^5}\) square units then you'll have \(\Large x*\sqrt{r^5}\) as the total area

jimthompson5910 (jim_thompson5910):

So therefore, \[\Large x*\sqrt{r^5} = 120r^3\] solve for x to get ???

OpenStudy (anonymous):

Okay so I'm confused on how to solve for x

jimthompson5910 (jim_thompson5910):

you divide both sides by sqrt(r^5), then rationalize the denominator like this \[\Large x*\sqrt{r^5} = 120r^3\] \[\Large x = \frac{120r^3}{\sqrt{r^5}}\] \[\Large x = \frac{120r^3*\sqrt{r^5}}{\sqrt{r^5}*\sqrt{r^5}}\] \[\Large x = \frac{120r^3*\sqrt{r^5}}{\sqrt{(r^5)^2}}\] \[\Large x = \frac{120r^3*\sqrt{r^5}}{r^5}\] \[\Large x = \frac{120r^3*\sqrt{r^{4+1}}}{r^5}\] \[\Large x = \frac{120r^3*\sqrt{r^{4}*r^{1}}}{r^5}\] \[\Large x = \frac{120r^3*\sqrt{r^{4}}*\sqrt{r}}{r^5}\] \[\Large x = \frac{120r^3*r^{2}*\sqrt{r}}{r^5}\] \[\Large x = \frac{120r^{3+2}*\sqrt{r}}{r^5}\] \[\Large x = \frac{120r^{5}*\sqrt{r}}{r^5}\] \[\Large x = 120\sqrt{r}\]

OpenStudy (anonymous):

Okay...so would the answer be that equation or what? I'm awful at this stuff

jimthompson5910 (jim_thompson5910):

well they want an expression, ie this \(\Large 120\sqrt{r}\) because that represents the exact number of pieces needed

jimthompson5910 (jim_thompson5910):

because we said "let x = number of pieces needed"

OpenStudy (anonymous):

So is the answer 120 sqrt. r?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

and those steps show why/how

OpenStudy (anonymous):

Okay so if i write all of that above and then make my final answer 120 sqrt. r thats it?

jimthompson5910 (jim_thompson5910):

yes pretty much

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome